# Privacy-Preserving Retrieval Augmented Generation with Differential Privacy

### v1

@@ Abstract
With the recent remarkable advancement of large language models (LLMs), there has been a growing interest in utilizing them in the domains with highly sensitive data that lies outside their training data. For this purpose, retrieval augmented generation (RAG) is particularly effective—it assists LLMs by directly providing relevant information from the external knowledge sources. However, without extra privacy safeguards, RAG outputs risk leaking sensitive information from the external data source. In this work, we explore RAG under differential privacy (DP), a formal guarantee of data privacy. The main challenge with differentially private RAG is how to generate long accurate answers within a moderate privacy budget. We address this by proposing an algorithm that smartly spends privacy budget only for the tokens that require the sensitive information and uses the non-private LLM for other tokens. Our extensive empirical evaluations reveal that our algorithm outperforms the non-RAG baseline under a reasonable privacy budget of $\epsilon\approx 10$ across different models and datasets.
@@ 1 Introduction
Large language models (LLMs) have shown a great deal of promise in a variety of applications. In particular, a major application of LLMs is in question-answering. The practical adoption of these systems often involves domains whose data is highly sensitive. For instance, healthcare institutions might want to utilize their internal medical records to provide precise medical information and personal feedback, while legal firms can leverage their case archives to assist clients with legal research and documentation. One way to achieve such domain-specific question-answering is through retrieval augmented generation (RAG) [5, 23, 16]. Here, we have a set of domain-specific documents; while answering a question, RAG retrieves a list of relevant documents and inputs them to LLMs as the context. However, even though this is effective for question-answering, RAG on a sensitive corpus can still leak private information about individual documents in this corpus [39, 30, 4, 29]. This is particularly problematic when end users are outside the data-holding entity, such as patients interacting with a healthcare institution’s RAG system.
@@
There are two aspects of challenges with designing an effective RAG algorithm under DP. The first is how to fit differential privacy into the RAG framework, and the second is how to manage the privacy-utility tradeoffs. We address the first challenge by proposing an algorithm, DPVoteRAG, based on the sample-and-aggregate framework in DP [26]. Our algorithm prepares multiple LLM instances, or voters, feeds disjoint partitions of the sensitive corpus to them, and produces output tokens one by one each through the majority vote of the voters’ token outputs. Note however that LLMs often output many tokens in response to a question. This is detrimental to privacy—the composition property of differential privacy states that multiple calculations based on the same dataset leads to higher privacy degradation. To resolve this challenge, we design another algorithm, DPSparseVoteRAG, that spends a privacy budget only when we need to. More specifically, we take advantage of the fact in RAG that LLMs require sensitive corpus only when generating tokens related to the knowledge. When not, outputs from LLMs without any context suffice. We formalize this idea with the sparse vector technique in DP [11, 12]—when voters agree with the non-private output of the LLM without contexts, we will simply output the non-private one without incurring a privacy budget. Consequently, our algorithm successfully generates sufficiently long, accurate responses under a reasonable privacy budget.
@@ 2 Preliminaries & Problem Setting
2.1 Retrieval Augmented Generation with Large Language Model
Retrieval augmented generation (RAG) is a technique to improve the performance of large language models (LLMs) on knowledge-intensive tasks by providing external knowledge. Given a question prompt, a retriever finds relevant documents from the external data source. Then, the relevant documents are added to the prompt as the contexts. An LLM (or generator) takes the augmented prompt as input and outputs the answer.
More formally, let $x\in\bigcup_{t=1}^{\infty}\mathcal{V}^{t}$ be a prompt, where $\mathcal{V}$ is some vocabulary. We further let $D$ be a dataset of documents as an external corpus with size $|D|=n$. A retriever $R$ finds a subset of $D$, $D_{x}\subset D$ that are relevant to $x$, i.e., $D_{x}=R(x,D)$. Finally, an LLM generates an answer $y=\mathrm{LLM}(x,D_{x})\in\bigcup_{t=1}^{\infty}\mathcal{V}^{t}$. The answer generation can further be decomposed into next-token generation. In particular, for each $t$, the $t$-th token $y_{t}\in\mathcal{V}$ is generated by $\mathrm{LLM}_{t}$, which takes $x$, $D_{x}$, and previously generated tokens $y_{<t}$ as inputs: $y_{t}=\mathrm{LLM}_{t}(x,D_{x},y_{<t})$.
@@ 2.2 Differential Privacy
Differential privacy is a strong cryptographically motivated definition of individual-level privacy. It guarantees that the participation of a single individual in a dataset does not change the probability of any outcome by much. In particular, suppose we have two datasets $D$ and $D^{\prime}$, each consisting of private data from $n$ individuals. We say that $D$ and $D^{\prime}$ are neighboring if they differ in a single individual’s private data. A randomized algorithm satisfies differential privacy if the output distributions on any pair of neighboring datasets are close. The formal definition is given as follows.
@@
One of the key properties of differential privacy is composition—sequential runs of differentially private algorithms also satisfy differential privacy. The composition property quantitatively captures the intuition that the more we release the information about the sensitive data, the worse the privacy guarantee becomes. More specifically, suppose $M_{1},\ldots,M_{T}$ are $(\epsilon_{0},\delta_{0})$-differentially private algorithms, which can be chosen adaptively based on previous outputs. Sequential composition theorem [9] states that the composed sequence of such algorithms guarantee $(T\epsilon_{0},T\delta_{0})$-differential privacy. Furthermore, advanced composition theorem [13, 19] states that the total privacy guarantee has $\epsilon=\mathcal{O}(\sqrt{T}\epsilon_{0})$.
@@ 2.2.1 Sparse Vector Technique
The sparse vector technique [11, 12] has originally emerged as the alternative of the composition in DP when we have so large number of numerical queries that the composition theorem of DP cannot provide a reasonable privacy guarantee but we are only interested in answers above some threshold. In such a case, the sparse vector technique algorithm, Sparse, reports whether each (noisy) query answer exceeds the threshold. It is shown that the privacy guarantee degrades by the number of queries above the threshold, instead of the total number of queries. Therefore, we save privacy budget by much when we expect only a few queries will be above the threshold. One can also report the (noisy) query outputs above the threshold by compromising privacy budget—the variant is called NumericSparse. The privacy guarantee follows by the composition of Sparse and differentially private query answering.
@@ 2.2.2 Differentially Private Generation via Sample-and-Aggregate
There has been a body of work on generating a token sequence by LLM with differential privacy. One common way to is to borrow the idea of the sample-and-aggregate framework in DP [26]. To generate a single token, a set of LLMs, each depending on a disjoint subset of the sensitive dataset $D$, generates a token or a probability distribution of a token respectively. When a set of tokens is generated, it forms an aggregate histogram of tokens, which is then carefully randomized with noise and only the most frequent token in the noisy histogram is published as the final output. When a set of probability distribution is generated, it forms a new mixture of probability distribution satisfying differential privacy. The final output token is then sampled from the resulting mixture distribution. The repetition of this process along with the composition theorem of DP yields the differentially private token sequence generation.
@@ 2.3 Problem Setting
Our goal is to generate an LLM answer to a prompt $x$ with retrieved external knowledge, $D_{x}=R(x,D)$, from a sensitive data source $D$ with differential privacy guarantee. More specifically, let the sensitive data source $D$ be a collection of individuals’ records—one record corresponds to one individual’s sensitive data [^1] It is straightforward to extend the setting to where multiple records correspond to one individual’s data by modifying the granularity of neighboring datasets in DP possibly with overhead in privacy-utility tradeoff.. We consider a realistic adversary who does not have direct access to the data source $D$ but have a capability of querying to the RAG system with any prompt $x$. We further assume that the LLM used in the RAG system is a copy of publicly available LLMs and is already pre-trained (and fine-tuned) with data disjoint from the sensitive data source $D$. That is, having access to the LLM parameters and/or pre-training (and fine-tuning) data does not provide any information on the sensitive data source $D$. To this end, we aim to formally guarantee that given any question $x$, a randomized LLM generation algorithm with RAG, $\mathrm{LLM}_{\mathrm{priv}}(x,R(x,D))$ satisfies $(\epsilon,\delta)$-differential privacy w.r.t the external knowledge data source $D$. We present the figure for this problem setting in Figure 1.
3 Differentially Private Retrieval Augmented Generation with Sparse Vector Technique
Figure 2: Overview of DPSparseVoteRAG when generating $t$-th token. It takes a question $x$, an interim answer up to $t-1$-th tokens, $y_{<t}$, and sensitive external knowledge $D$ as inputs and outputs a new token $y_{t}$ or $y_{t}^{\mathrm{non-RAG}}$. Components enclosed in red indicate that operations are done in a privacy-preserving manner. DPVoteRAG works the same until computing the output token histogram $h_{t}$, but it immediately finds the most frequent element in $h_{t}$ afterwards. Note that each retriever output only depends on $x$ and the corresponding subset of $D$; thus, the algorithm needs not to repeat the retrieval process for every token generation. We add the retrieval procedure in the figure for completeness.
Our differentially private RAG algorithm consists of two main components—DP voting for the single-token generation and efficient privacy budget spending by leveraging the sparse vector technique combined with the utilization of LLMs without any relevant documents provided. These two components enable us to generate answers that incorporate external knowledge while guaranteeing reasonable level of differential privacy. We start from our algorithm with the first component alone, and then extend it to include the second component. The graphical overview of our algorithm is presented in Figure 2.
@@ 3.1 DPVoteRAG: Differentially Private Voting Algorithm for RAG
Input: Prompt $x$, External data source $D$, Generator $\mathrm{LLM}$, Retriever $R$, # of voters $m$, # of retrieval per voter $k$, Per-token privacy budget $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$, Total privacy budget $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$
Output: Private answer $y$
1
@@
3 $D_{1},\ldots,D_{m}\leftarrow$ split $D$ into $m$ disjoint subset;
4
5for $t\leftarrow 1$ to $T_{\max}$ do
6       for $i\leftarrow 1$ to $m$ do
             $y_{t}^{(i)}\leftarrow\mathrm{LLM}_{t}(x,R(x,D_{i}),y_{<t})$;
              // Generate a token with RAG
7
8       end for
9      $h_{t}\leftarrow$ Build a histogram of tokens from $y_{t}^{(1)},\ldots,y_{t}^{(m)}$;
       $y_{t}\leftarrow\mathrm{LimitedDomain}(h_{t},\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$;
        // Privately choose the most frequent token
[…]
             return $(y_{1},\ldots,y_{t})$;
              // Halt if end of sequence
11
12       end if
13
14 end for
15return $(y_{1},\ldots,y_{T_{\max}})$;
16
Algorithm 1 DPVoteRAG
By the nature of retrieval in RAG—retrieving relevant documents for a question, the LLM outputs can depend on a sensitive individual’s document. Therefore, our algorithmic design needs to relax the dependency of a single individual’s document on the output, while exploiting the external data source, to achieve a reasonable privacy-utility tradeoff. Inspired by the differentially private generation via sample-and-aggregate framework in Section 2.2.2, we present a differentially private voting algorithm for RAG—DPVoteRAG.
Given a prompt $x$ and external data source $D$, DPVoteRAG first partitions $D$ into $m$ disjoint datasets $D_{1},\ldots,D_{m}$. Then, for each $i=1,\ldots,m$, it retrieves $k$ relevant documents from $D_{i}$, feeds them into the LLM along with the original prompt $x$, and generates a next token. It collects these tokens to form a histogram over the vocabulary. It remains to privately choose the most frequent element from the histogram. While it is generally hard to do so when the histogram dimension is large as in our setting, e.g., the vocabulary size of OPT [41] is 50272, there is a line of work in the community to overcome this difficulty. Following the work by [15], we integrate the LimitedDomain mechanism [8] into our algorithm. The mechanism enables us to reduce the histogram dimension significantly with some cost in privacy budget and thus achieve a better privacy-utility tradeoff. By its design, the LimitedDomain mechanism possibly outputs the designed null token. In such a case, we halt the algorithm, or equivalently, regard that it outputs the end of sequence token.  [^2] We find that by choosing the appropriate size of reduced dimension, the LimitedDomain mechanism in our experiment rarely outputs the null token. Finally, we append the chosen token to the next input to the LLM. We repeat this process until we see the end of sequence token chosen or reach the maximum number of token generation, which is computed in advance from the per-token and total privacy budget  [^3] The maximum number of token generation is computed as follows. We first calculate the maximum numbers of composition with the sequential and advanced composition theorem [12] under the per-token privacy budget $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$ and total privacy budget $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$. Then, we take the maximum of two numbers of possible composition.. We present the concrete algorithm in Algorithm 1. The formal privacy analysis is as follows.
@@ 9      $h_{t}\leftarrow$ Build a histogram of tokens from $y_{t}^{(1)},\ldots,y_{t}^{(m)}$;
The guarantee simply follows from the privacy guarantee of the LimitedDomain mechanism and the composition theorem used to compute $T_{\max}$ in the algorithm.

### v2

# Privacy-Preserving Retrieval-Augmented Generation with Differential Privacy

@@ Abstract
With the recent remarkable advancement of large language models (LLMs), there has been a growing interest in utilizing them in the domains with highly sensitive data that lies outside their training data. For this purpose, retrieval-augmented generation (RAG) is particularly effective—it assists LLMs by directly providing relevant information from the external knowledge sources. However, without extra privacy safeguards, RAG outputs risk leaking sensitive information from the external data source. In this work, we explore RAG under differential privacy (DP), a formal guarantee of data privacy. The main challenge with differentially private RAG is how to generate long accurate answers within a moderate privacy budget. We address this by proposing an algorithm that smartly spends privacy budget only for the tokens that require the sensitive information and uses the non-private LLM for other tokens. Our extensive empirical evaluations reveal that our algorithm outperforms the non-RAG baseline under a reasonable privacy budget of $\epsilon\approx 10$ across different models and datasets.
@@ 1 Introduction
Large language models (LLMs) have shown a great deal of promise in a variety of applications. In particular, a major application of LLMs is in question-answering (QA). The practical adoption of these systems often involves domains whose data is highly sensitive. For instance, healthcare institutions might want to utilize their internal medical records to provide precise medical information and personal feedback, while legal firms can leverage their case archives to assist clients with legal research and documentation. One way to achieve such domain-specific QA is through retrieval-augmented generation (RAG) [5, 24, 17]. Here, we have a set of domain-specific documents; while answering a question, RAG retrieves a list of relevant documents and inputs them to LLMs as the context. However, even though this is effective for QA, RAG on a sensitive corpus can leak private information about individual documents in the corpus [40, 31, 4, 30]. This is particularly problematic when end users are outside the data-holding entity, e.g., patients interacting with a healthcare institution’s RAG system.
@@
There are two aspects of the challenges with designing an effective RAG algorithm under DP. The first is how to fit differential privacy into the RAG framework, and the second is how to manage the privacy-utility tradeoffs. We address the first challenge by proposing an algorithm, DPVoteRAG, based on the sample-and-aggregate framework in DP [27]. Our algorithm prepares multiple LLM instances, or voters, feeds disjoint partitions of the sensitive corpus to them, and produces output tokens one by one each through the majority vote of the voters’ token outputs. Note, however, that LLMs often output many tokens in response to a question. This is detrimental to privacy—the composition property of differential privacy states that multiple calculations based on the same dataset lead to greater privacy degradation. To resolve this challenge, we design another algorithm, DPSparseVoteRAG, that spends a privacy budget only when we need to. More specifically, we take advantage of the fact in RAG that LLMs require the sensitive corpus only when generating tokens related to the knowledge. When not, outputs from LLMs without any context suffice. We formalize this idea with the sparse vector technique in DP [12, 13]—when voters agree with the non-private output of the LLM without contexts, we will simply output the non-private one without incurring a privacy budget. Consequently, our algorithm successfully generates sufficiently long, accurate responses under a reasonable privacy budget.
@@ 2 Preliminaries & Problem Setting
2.1 Retrieval-Augmented Generation with Large Language Model
Retrieval-augmented generation (RAG) is a technique to improve the performance of large language models (LLMs) on knowledge-intensive tasks by providing external knowledge. Given a question prompt, a retriever finds relevant documents from the external data source. Then, the relevant documents are added to the prompt as the contexts. An LLM (or generator) takes the augmented prompt as input and outputs the answer.
More formally, let $x\in\bigcup_{t=1}^{\infty}\mathcal{V}^{t}$ be a prompt, where $\mathcal{V}$ is some vocabulary. We further let $D$ be a dataset of documents as an external corpus with size $|D|=n$. A retriever $R$ finds a subset of $D$, $D_{x}\subset D$, with size $k$ that is relevant to $x$, i.e., $D_{x}=R(x,D;k)$. Finally, an LLM generates an answer $y=\mathrm{LLM}(x,D_{x})\in\bigcup_{t=1}^{\infty}\mathcal{V}^{t}$. The answer generation can further be decomposed into next-token generation. In particular, for each $t$, the $t$-th token $y_{t}\in\mathcal{V}$ is generated by $\mathrm{LLM}_{t}$, which takes $x$, $D_{x}$, and previously generated tokens $y_{<t}$ as inputs: $y_{t}=\mathrm{LLM}_{t}(x,D_{x},y_{<t})$.
@@ 2.2 Differential Privacy
Differential privacy (DP) is a strong cryptographically motivated definition of individual-level privacy. It guarantees that the participation of a single individual in a dataset does not change the probability of any outcome by much. In particular, suppose we have two datasets $D$ and $D^{\prime}$, each consisting of private data from $n$ individuals. We say that $D$ and $D^{\prime}$ are neighboring if they differ in a single individual’s private data. A randomized algorithm satisfies differential privacy if the output distributions on any pair of neighboring datasets are close. The formal definition is given as follows.
@@
One of the key properties of DP is composition—sequential runs of differentially private algorithms also satisfy differential privacy. The composition property quantitatively captures the intuition that the more we release the information about the sensitive data, the worse the privacy guarantee becomes. More specifically, suppose $M_{1},\ldots,M_{T}$ are $(\epsilon_{0},\delta_{0})$-differentially private algorithms, which can be chosen adaptively based on previous outputs. Sequential composition theorem [10] states that the composed sequence of such algorithms guarantee $(T\epsilon_{0},T\delta_{0})$-differential privacy. Furthermore, advanced composition theorem [14, 20] states that the total privacy guarantee has $\epsilon=\mathcal{O}(\sqrt{T}\epsilon_{0})$.
@@ 2.2.1 Sparse Vector Technique
The sparse vector technique [12, 13] has originally emerged as the alternative of the composition in DP when we have such a large number of numerical queries that the composition theorem cannot provide a reasonable privacy guarantee but we are only interested in answers above some threshold. In such a case, the sparse vector technique algorithm, Sparse, reports whether each (noisy) query answer exceeds the threshold. It is shown that the privacy guarantee degrades by the number of queries above the threshold, instead of the total number of queries. Therefore, we save privacy budget by much when we expect only a few queries will be above the threshold. One can also report the (noisy) query outputs above the threshold by compromising privacy budget—the variant is called NumericSparse. The privacy guarantee follows by the composition of Sparse and differentially private query answering.
@@ 2.2.2 Differentially Private Generation via Sample-and-Aggregate
There has been a body of work on generating a token sequence by LLM with DP. The most common way is to borrow the idea of the sample-and-aggregate framework in DP [27]. To generate a single token, a set of LLMs, each depending on a disjoint subset of the sensitive dataset $D$, generates a token respectively. The generated tokens form an aggregate histogram of tokens, which is then carefully randomized with noise and only the most frequent token in the noisy histogram is published as the final output. The repetition of this process along with the composition theorem of DP yields the differentially private token sequence generation.
@@ 2.3 Problem Setting
Our goal is to generate an LLM answer to a prompt $x$ with retrieved external knowledge, $D_{x}=R(x,D;k)$, from a sensitive data source $D$ with a differential privacy guarantee. More specifically, let the sensitive data source $D$ be a collection of individuals’ records—one record corresponds to one individual’s sensitive data [^1] It is straightforward to extend the setting to where multiple records correspond to one individual’s data by modifying the granularity of neighboring datasets in DP possibly with overhead in privacy-utility tradeoff.. We consider a realistic adversary who does not have direct access to the data source $D$ but has a capability of querying to the RAG system with any prompt $x$. We further assume that the LLM used in the RAG system is a copy of publicly available LLMs and is already pre-trained (and fine-tuned) with data disjoint from the sensitive data source $D$. That is, having access to the LLM parameters and/or pre-training (and fine-tuning) data does not provide any information on the sensitive data source $D$. To this end, we aim to formally guarantee that given any question $x$, a randomized LLM generation algorithm with RAG, $\mathrm{LLM}_{\mathrm{priv}}(x,R(x,D;k))$ satisfies $(\epsilon,\delta)$-differential privacy w.r.t the external knowledge data source $D$. We present the figure for this problem setting in Figure 1.
3 Differentially Private Retrieval-Augmented Generation with Sparse Vector Technique
Figure 2: Overview of DPSparseVoteRAG. It first retrieves the relevant documents to a question $x$ from sensitive external knowledge $D$ and partitions them uniform randomly (left). When generating $t$-th token (right), it takes a question $x$ and an interim answer up to $t-1$-th tokens, $y_{<t}$, as inputs and outputs a new token $y_{t}$. Components enclosed in red indicate that operations involve randomness for the privacy guarantee. DPVoteRAG works the same until computing the output token histogram $h_{t}$, but it immediately finds the majority token in $h_{t}$ afterwards.
Our differentially private RAG algorithm consists of two main components—DP voting for the single-token generation and efficient privacy budget spending by leveraging the sparse vector technique combined with the utilization of LLMs without any relevant documents provided. These two components enable us to generate answers that incorporate external knowledge while guaranteeing a reasonable level of differential privacy. We start from our algorithm with the first component alone, and then extend it to include the second component. The graphical overview of our algorithm is presented in Figure 2.
@@ 3.1 DPVoteRAG: Differentially Private Voting Algorithm for RAG
Algorithm 1 DPVoteRAG
0:  Prompt $x$, External data source $D$, Generator $\mathrm{LLM}$, Retriever $R$, # of voters $m$, # of retrieval per voter $k$, Per-token privacy budget $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$, Total privacy budget $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$
0:  Private answer $y$
1:  $T_{\max}\leftarrow$ maximum # of tokens to generate based on $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$ and $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$
2:  // Retrieval and random partition of the relevant documents
3:  $D_{x}\leftarrow$ Retrieve $mk$ most relevant documents $R(x,D;mk)$.
4:  $D_{x}^{1},\cdots,D_{x}^{m}\leftarrow$ Uniformly randomly partition $D_{x}$ into $m$ disjoint subsets.
5:  for $t\leftarrow 1$ to $T_{\max}$ do
6:     // Generating the token histogram with the sensitive documents
7:     for $i\leftarrow 1$ to $m$ do
8:        $y_{t}^{(i)}\leftarrow\mathrm{LLM}_{t}(x,D_{x}^{i},y_{<t})$
9:     end for
10:     $h_{t}\leftarrow$ Build a histogram of tokens from $y_{t}^{(1)},\ldots,y_{t}^{(m)}$
11:     // Producing a token from the histogram privately
12:     $y_{t}\leftarrow\mathrm{LimitedDomain}(h_{t},\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$
13:     // Halting if end of sequence
@@
15:        return $(y_{1},\ldots,y_{t})$
16:     end if
17:  end for
18:  return $(y_{1},\ldots,y_{T_{\max}})$
By the nature of retrieval in RAG—retrieving relevant documents for a question, the LLM outputs can depend on a sensitive individual’s document. Therefore, our algorithmic design needs to relax the dependency of a single individual’s document on the output, while exploiting the external data source, to achieve a reasonable privacy-utility tradeoff. Inspired by the differentially private generation via sample-and-aggregate framework, we present a differentially private voting algorithm for RAG—DPVoteRAG.
Given a prompt $x$ and external data source $D$, DPVoteRAG first retrieves $mk$ documents as $D_{x}$. Then it makes uniformly randomly partitions $D_{x}$ into $m$ disjoint datasets $D_{x}^{1},\ldots,D_{x}^{m}$ and each subset has exact size $k$. Then, for each $i=1,\ldots,m$, it feeds $k$ documents $D_{x}^{i}$ into the LLM along with the original prompt $x$, and generates a next token. It collects these tokens to form a histogram over the vocabulary. It remains to privately choose the most frequent element from the histogram. While it is generally hard to do so when the histogram dimension is large as in our setting, e.g., the vocabulary size of OPT [42] is 50272, there is a line of work in the community to overcome this difficulty. Following the work by [16], we integrate the LimitedDomain mechanism [9] into our algorithm. The mechanism enables us to reduce the histogram dimension significantly with some cost in a privacy budget and thus achieve a better privacy-utility tradeoff. By its design, the LimitedDomain mechanism possibly outputs the designed null token. In such a case, we halt the algorithm, or equivalently, regard that it outputs the end of sequence token.  [^2] We find that by choosing the appropriate size of reduced dimension, the LimitedDomain mechanism in our experiment rarely outputs the null token. Finally, we append the chosen token to the next input to the LLM. We repeat this process until we see the end of sequence token chosen or reach the maximum number of token generation, which is computed in advance from the per-token and total privacy budget  [^3] The maximum number of token generation is computed as follows. We first calculate the maximum numbers of composition with the sequential and advanced composition theorem [13] under the per-token privacy budget $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$ and total privacy budget $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$. Then, we take the maximum of two numbers of possible composition.. We present the concrete algorithm in Algorithm 1. The formal privacy analysis is as follows.
@@
The guarantee simply follows from the property of uniformly random partition, the privacy guarantee of the LimitedDomain mechanism and the composition theorem used to compute $T_{\max}$. The full proof is in Appendix A.